OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 343 cm3?

OpenStudy (anonymous):

Volume of a cube = Length^3 So Length = Volume^(1/3)

OpenStudy (anonymous):

so i divide 343 by 3?

OpenStudy (anonymous):

nope thats actually a cubic root \[\sqrt[3]{x}=x^{1/3}\]

OpenStudy (whpalmer4):

\[v=l^3\]\[343=l^3\]\[\sqrt[3]{343}=l\]\[l=\]

OpenStudy (anonymous):

thankyou

OpenStudy (whpalmer4):

In other words, \[l*l*l=343\] You can find it by trying smallish integers.

OpenStudy (whpalmer4):

5*5*5=125 10*10*10=1000 Must be between them